Grade 12Physical Chemistry

A gas absorbs a photon of 355 nm and emits at two wavelengths. If one of the emission is at 680nm the other is at

Alexzander
9 Years agoGrade 12

1 Answer

Vikas TU
9 Years ago
energy corresponding to wavelength 355nm=hc/lwhere h= planks constantc=speed of lightl=wavelengththerefore according to the problem,hc/355*10^-9=hc/680^10^-9 +hc/l*10^-91/355=1/680+1/ll=742nmThus the wavelength of the light for other emmission is 742nm.